September 30, 2022
- Switch overed from CaFe to XEM2 experiments.
- SHMS/HMS trigger setups were checked/modified according to the XEM2 experiment trigger requirements https://logbooks.jlab.org/entry/4048532, https://logbooks.jlab.org/entry/4048593
- Data file locations were modified for the XEM2 experiments as well https://logbooks.jlab.org/entry/4048556.
- Beam was restored to the Hall C around 22:45. We did a quick harp scan and beam centering check with the Al hole target while we had beam briefly https://logbooks.jlab.org/entry/4048797 https://logbooks.jlab.org/entry/4048844
- Beam was interrupted due to the 1L12 trip and all RF in the North Linac. This was followed by the RSEP8A magnet over temperature fault. DC power is looking into it.
- HMS Q1, Q2 interlocks twice while ramping up to the MOL. Was able to set the magnets up by increasing the current in increments.
- Problem with the HMS rotation to 19.52 deg. Engineer on call came in to reset the drive controller. Worked at the second attempt.
- HMS Q3 and Dipole interlocks during the early owl shift - reset worked. https://logbooks.jlab.org/entry/4048848
- Due to the older hcana version on our go_analysis, we were not seeing any events on the fast raster plots. Casey switched to the newer firmware version - problem is fixed.
Mark also confirmed that with the new FADC firmware, we need to use the latest HCANA.
- Nadia noticed the HMS Q1 LN2 levels have been dropping since last night. https://logbooks.jlab.org/entry/4048919 Seems recovering https://logbooks.jlab.org/entry/4048927
Things need to be fixed:
- Prescale GUI problem https://logbooks.jlab.org/entry/4048783
- Ezekiel Wertz came to notify us that one of the two ethane bottles in the gas shed is empty. It needs to be replayed at some time soon.